# Casing Collar Identification using AlexNet-based Neural Networks for Depth Measurement in Oil and Gas Wells

### v1

@@ Abstract
Accurate downhole depth measurement is essential for oil and gas well operations, directly influencing reservoir contact, production efficiency, and operational safety. Collar correlation using a casing collar locator (CCL) is fundamental for precise depth calibration. While neural network-based CCL signal recognition has achieved significant progress in collar identification, preprocessing methods for such applications remain underdeveloped. Moreover, the limited availability of real well data poses substantial challenges for training neural network models that require extensive datasets. This paper presents a system integrated into downhole tools for CCL signal acquisition to facilitate dataset construction. We propose comprehensive preprocessing methods for data augmentation and evaluate their effectiveness using our AlexNet-based neural network models. Through systematic experimentation across various configuration combinations, we analyze the contribution of each augmentation method. Results demonstrate that standardization, label distribution smoothing (LDS), and random cropping are fundamental requirements for model training, while label smoothing regularization (LSR), time scaling, and multiple sampling significantly enhance model generalization capability. The F1 scores of our two benchmark models trained with the proposed augmentation methods maximumly improve from 0.937 and 0.952 to 1.0 and 1.0, respectively. Performance validation on real CCL waveforms confirms the effectiveness and practical applicability of our approach. This work addresses the gaps in data augmentation methodologies for training casing collar recognition models in CCL data-limited environments.
@@ I Introduction
Figure 1: Cross-sectional illustration of a typical oil and gas well structure. Representative casing collar signals from magnetic response are illustrated in dark blue near the corresponding casing collar, while the typical interference signals distant from casing collars are illustrated in dark green. Refer to figures from [1].
@@
Although a length measuring wheel (LMW) is able to estimate depth as tools descend during wireline intervention operations, its accuracy remains insufficient. To improve depth measurement accuracy, calibration using signals from a casing collar locator (CCL) – a magnetic downhole positioning tool that detects magnetic anomalies at the casing collars of wellbores or pipes – is fundamental [4]. The CCL produces a characteristic magnetic response as an electrical signal when passing through each collar, termed a “collar signal” or “collar magnetic signature”. The characteristic magnetic response pattern typically exhibits a bimodal distribution, as illustrated by dark blue waveforms in Fig. 1. Through collar correlation, which refers to matching CCL logs with casing tally, depth reference markers are established, enabling actual and accurate depth measurement [5, 6]. The casing tally, also known as the list of collars or casing string reference depths, records the depths of casing collars and is commonly extracted from cementing quality data.
While collar correlation using CCL signals is an established method, collar signal recognition presents significant challenges. CCL signal integrity can be severely compromised by multiple factors, including cable effects, wellbore conditions [7], tool motion (swing or rotation), amplifier saturation, and environmental noise [5, 8]. Consequently, collar signal waveforms become increasingly ambiguous, necessitating robust recognition of collar signals amid various interferences, as illustrated by dark green waveforms in Fig. 1.
Various signal processing techniques have been developed to identify collar signals under interference conditions. Traditional methods include fixed or dynamic thresholding [1, 9], digital filters and template-based cross-correlation [10], time-frequency domain techniques such as Fourier or wavelet transforms [11], and physical plausibility filters [1, 4, 12]. However, these approaches exhibit limited generalizability [8, 12]. With the emergence of machine learning, researchers have increasingly employed deep neural networks to automate collar signal recognition. These developments include convolutional neural networks (CNNs) and long short-term memory networks (LSTMs) [13, 14, 15], Additionally, advanced architectures such as transformer models [16] and physics-informed neural networks (PINNs) [17] have been proposed for downhole signal classification, anomaly detection, and denoising tasks [18].
@@
The limited sample amount and unique characteristics of CCL signal data necessitate specialized preprocessing methods for neural network training. However, research on preprocessing methods remain scarce, despite extensive work on collar signal recognition [4, 8, 10, 11, 12, 13, 16, 18, 17]. Fortunately, extensive work exists on data augmentation for preventing overfitting and improving generalizability when training on small datasets. Data normalization methods – including min-max scaling, Z-score normalization (standardization), and robust scaling – demonstrate important roles in stabilizing input distributions and gradients, preventing vanishing and exploding gradients, and ultimately accelerating training convergence [20, 21, 22, 23]. Label smoothing regularization (LSR) and label distribution smoothing (LDS) discourage overconfident predictions, thereby enhancing model generalizability [24, 25]. Similarly, employing probability maps for boundary prediction instead of one-hot encoding (OHE) labels through boundary probabilization transforms the training objective from single-point prediction to probability distribution estimation, enabling models learn smooth and fuzzy decision boundaries that are more resistant to noise or perturbations [24, 14, 26]. Notably, Gaussian kernel often deliver optimal results in LDS applications [24]. Furthermore, graph augmentation techniques including randomly cropping, scaling, translation, and noise injection expand training datasets while improving model generalizability and robustness [27, 28].
@@
  • We propose two neural networks models for collar signal recognition that serve as benchmarks for evaluating data preprocessing methods, as illustrated in Fig. 3. The first model, Thin AlexNet (TAN), is modified from AlexNet – a classic and proven architecture in pattern recognition. The second model, Miniaturized AlexNet (MAN), is a simplified version of TAN with fewer layers.
@@ II-A Problem Transformation
As previously discussed, accurately positioning downhole tools through collar correlation relies on correct identification of bimodal patterns (i.e., collar signals) within CCL signals. The center of each bimodal pattern is typically considered as the moment when the CCL coincides with the casing collar.
By observing signal characteristics, collar signals can be clearly identified from signals proximate to the corresponding collar rather than from distant signals. Therefore, collar signal identification is largely independent of CCL signals on large time scales. Consequently, focusing on CCL signal fragments using an appropriately sized sliding window is reasonable, consistent with approaches in [14] and [15]. When CCL signals are sampled at a fixed frequency, the absolute value of timestamps within sliding windows becomes irrelevant, indicating that the pure CCL signal sequence sufficiently conveys the necessary information.
As shown in Fig. 1, the casing tally indicates a series of depths that can be mapped to collar signals on the signal timeline as a series of marks. The center points of collar signals marked on the timeline (hereafter referred to as “collar marks”) constitute a series of timestamps that require different representation when captured by sliding windows. The most direct method for representing collars marks is one-hot encoding (OHE), where 1 denotes “is a collar mark” and 0 denotes “is not a collar mark”. This approach transforms collar mark prediction into a binary classification problem.
@@ III-A Evaluation Measures
To recognize collar signals from normalized long CCL waveforms, a sliding window of width $W$ with a strid of $\lfloor W/2\rfloor$ is employed. This configuration produces 50% overlap between consecutive windows, ensuring that each overlapping region is captured by exactly two adjacent windows. The window advances through the waveform in half-width increments, balancing computational efficiency with temporal resolution while maintaining analytical continuity. The probability map segments inferenced from overlapping regions of adjacent windows are averaged to produce the complete probability map, as illustrated in Fig. 3(f).

### v2

# Data-Augmented Deep Learning for Downhole Depth Sensing and Field Validation

@@ Abstract
Accurate downhole depth measurement is essential for oil and gas well operations, directly influencing reservoir contact, production efficiency, and operational safety. Collar correlation using a casing collar locator (CCL) is fundamental for precise depth calibration. While neural network-based CCL signal recognition has achieved significant progress in collar identification, preprocessing methods for such applications remain underdeveloped. Moreover, the limited availability of real well data poses substantial challenges for training neural network models that require extensive datasets. This paper presents a system integrated into downhole tools for CCL signal acquisition to facilitate dataset construction. We propose comprehensive preprocessing methods for data augmentation and evaluate their effectiveness using our neural network models. Through systematic experimentation across various configuration combinations, we analyze the contribution of each augmentation method. Results demonstrate that standardization, label distribution smoothing (LDS), and random cropping are fundamental requirements for model training, while label smoothing regularization (LSR), time scaling, and multiple sampling significantly enhance model generalization capability. The F1 scores of our two benchmark models trained with the proposed augmentation methods maximumly improve from 0.937 and 0.952 to 1.0 and 1.0, respectively. Performance validation on real CCL waveforms confirms the effectiveness and practical applicability of our approach. This work addresses the gaps in data augmentation methodologies for training casing collar recognition models in CCL data-limited environments.
@@ I Introduction
Figure 1: Cross-sectional illustration of a typical oil and gas well structure. Representative casing collar signatures from magnetic response are illustrated in dark blue near the corresponding casing collar, while the typical interference signals distant from casing collars are illustrated in dark green. Refer to figures from [1].
@@
Although a length measuring wheel (LMW) is able to estimate depth as tools descend during wireline intervention operations, its accuracy remains insufficient. To improve depth measurement accuracy, calibration using signals from a casing collar locator (CCL) – a magnetic downhole positioning tool that detects magnetic anomalies at the casing collars of wellbores or pipes – is fundamental [4]. The CCL produces a characteristic magnetic response as an electrical signal when passing through each collar [4, 5, 6], termed a “collar signal” or “collar (magnetic) signature”. The characteristic magnetic response pattern typically exhibits a bimodal distribution, as illustrated by dark blue waveforms in Fig. 1. Through collar correlation, which refers to matching CCL logs with casing tally, depth reference markers are established, enabling actual and accurate depth measurement [7, 6]. The casing tally, also known as the list of collars or casing string reference depths, records the depths of casing collars and is commonly extracted from cementing quality data.
While collar correlation using CCL signals is an established method, collar signature recognition presents significant challenges. CCL signal integrity can be severely compromised by multiple factors, including cable effects, wellbore conditions [8], tool motion (swing or rotation), amplifier saturation, and environmental noise [7, 9]. Consequently, collar signature waveforms become increasingly ambiguous, necessitating robust recognition of collar signatures amid various interferences, as illustrated by dark green waveforms in Fig. 1.
Various signal processing techniques have been developed to identify collar signatures under interference conditions. Traditional methods include fixed or dynamic thresholding [1, 10], digital filters and template-based cross-correlation [11], time-frequency domain techniques such as Fourier or wavelet transforms [12], and physical plausibility filters [1, 4]. However, these approaches exhibit limited generalizability [9, 13]. With the emergence of machine learning, researchers have increasingly employed deep neural networks to automate collar signature recognition. These developments include convolutional neural networks (CNNs) and long short-term memory networks (LSTMs) [14, 15, 16], Additionally, advanced architectures such as transformer models [17] and physics-informed neural networks (PINNs) [18] have been proposed for related fields such as downhole signal classification, anomaly detection, and denoising tasks [19, 20, 21, 22, 13].
@@
The limited sample amount and unique characteristics of CCL signal data necessitate specialized preprocessing methods for neural network training. However, research on preprocessing methods remain scarce, despite extensive work on collar signature recognition [4, 9, 11, 12, 13, 14, 17, 19, 18]. Fortunately, extensive work exists on data augmentation for preventing overfitting and improving generalizability when training on small datasets. Data normalization methods – including min-max scaling, Z-score normalization (standardization), and robust scaling – demonstrate important roles in stabilizing input distributions and gradients, preventing vanishing and exploding gradients, and ultimately accelerating training convergence [24, 25, 26, 27]. Label smoothing regularization (LSR) and label distribution smoothing (LDS) discourage overconfident predictions, thereby enhancing model generalizability [28, 29]. Similarly, employing probability maps for boundary prediction instead of one-hot encoding (OHE) labels through boundary probabilization transforms the training objective from single-point prediction to probability distribution estimation, enabling models learn smooth and fuzzy decision boundaries that are more resistant to noise or perturbations [28, 15, 30]. Notably, Gaussian kernel often deliver optimal results in LDS applications [28]. Furthermore, graph augmentation techniques including randomly cropping, scaling, translation, and noise injection expand training datasets while improving model generalizability and robustness [31, 32].
@@
  • We propose two neural networks models for collar signature recognition that serve as benchmarks for evaluating data preprocessing methods, as illustrated in Fig. 3. The first model, Thin AlexNet (TAN), is modified from AlexNet – a classic and proven architecture in pattern recognition. The second model, Miniaturized AlexNet (MAN), is a simplified version of TAN with fewer layers.
@@ II-A Problem Transformation
As previously discussed, accurately positioning downhole tools through collar correlation relies on correct identification of bimodal patterns (i.e., collar signatures) within CCL signals. The center of each bimodal pattern is typically considered as the moment when the CCL coincides with the casing collar.
By observing signal characteristics, collar signatures can be clearly identified from signals proximate to the corresponding collar rather than from distant signals. Therefore, collar signature recognition is largely independent of CCL signals on large time scales. Consequently, focusing on CCL signal fragments using an appropriately sized sliding window is reasonable, consistent with approaches in [15] and [16]. When CCL signals are sampled at a fixed frequency, the absolute value of timestamps within sliding windows becomes irrelevant, indicating that the pure CCL signal sequence sufficiently conveys the necessary information.
As shown in Fig. 1, the casing tally indicates a series of depths that can be mapped to collar signatures on the signal timeline as a series of marks. The center points of collar signatures marked on the timeline (hereafter referred to as “collar marks”) constitute a series of timestamps that require different representation when captured by sliding windows. The most direct method for representing collars marks is one-hot encoding (OHE), where 1 denotes “is a collar mark” and 0 denotes “is not a collar mark”. This approach transforms collar mark prediction into a binary classification problem.
@@ III-A Evaluation Measures
To recognize collar signatures from normalized long CCL waveforms, a sliding window of width $W$ with a strid of $\lfloor W/2\rfloor$ is employed. This configuration produces 50% overlap between consecutive windows, ensuring that each overlapping region is captured by exactly two adjacent windows. The window advances through the waveform in half-width increments, balancing computational efficiency with temporal resolution while maintaining analytical continuity. The probability map segments inferenced from overlapping regions of adjacent windows are averaged to produce the complete probability map, as illustrated in Fig. 3(f).